OpenStudy (anonymous):

f(x)=(e^x)*(5x^3-5x^2+2x)*(tan(x)) d/dx f(x)=?

OpenStudy (anonymous):

looks like a product rule problem to me yes?

OpenStudy (anonymous):

\[(fgh)'=f'gh+fg'h+fgh'\]

OpenStudy (anonymous):

so put \[f(x)=e^x,f'(x)=e^x\] \[g(x)=5x^3-5x^2+2x,g'(x)=15x^2-10x+2\] \[h(x)=\tan(x),h'(x)=\sec^2(x)\] and go to town. clear now or should i finish?

OpenStudy (anonymous):

thanks you

OpenStudy (anonymous):

any idea now, of what the derivative of ln (x) is??

OpenStudy (anonymous):

yes it is \[\frac{1}{x}\]